# Hybrid Selective State Models

### v1

@@ Abstract
Diffusion models have been widely employed in the field of 3D manipulation due to their efficient capability to learn distributions, allowing for precise prediction of action trajectories. However, diffusion models typically rely on large parameter UNet backbones as policy networks, which can be challenging to deploy on resource-constrained devices. Recently, the Mamba model has emerged as a promising solution for efficient modeling, offering low computational complexity and strong performance in sequence modeling. In this work, we propose the Mamba Policy, a lighter but stronger policy that reduces the parameter count by over 80% compared to the original policy network while achieving superior performance. Specifically, we introduce the XMamba Block, which effectively integrates input information with conditional features and leverages a combination of Mamba and Attention mechanisms for deep feature extraction. Extensive experiments demonstrate that the Mamba Policy excels on the Adroit, Dexart, and MetaWorld datasets, requiring significantly fewer computational resources. Additionally, we highlight the Mamba Policy’s enhanced robustness in long-horizon scenarios compared to baseline methods and explore the performance of various Mamba variants within the Mamba Policy framework. Our project page is in this link.
@@ IV-B XMamba
TABLE I: Quantitative Comparisons of different baselines in simulation environments. We compare our Mamba Policy with IBC, BCRNN, 3D Diffusion Policy, and Diffusion Policy in Adroit, MetaWorld and DexArt datasets. denotes our reproduced results for fair comparison. Our proposed Mamba Policy achieves superior results across all domains.
@@
3D Diffusion Policy [4] $100\scriptstyle{\pm 0}$ $62\scriptstyle{\pm 4}$ $43\scriptstyle{\pm 6}$ $99\scriptstyle{\pm 1}$ $69\scriptstyle{\pm 4}$ $97\scriptstyle{\pm 4}$ $83\scriptstyle{\pm 1}$ $63\scriptstyle{\pm 2}$ $82\scriptstyle{\pm 4}$ $46\scriptstyle{\pm 2}$ $74.4$
@@
DP3 $100.0\scriptstyle{\pm 0.0}$ $64.5\scriptstyle{\pm 2.4}$ $41.0\scriptstyle{\pm 3.3}$ $99.6\scriptstyle{\pm 0.4}$ $\mathbf{76.0\scriptstyle{\pm 2.4}}$ $\mathbf{100.0\scriptstyle{\pm 0.0}}$ $77.0\scriptstyle{\pm 3.5}$ $30.0\scriptstyle{\pm 2.9}$ $74.3\scriptstyle{\pm 2.6}$ $27.0\scriptstyle{\pm 0.8}$ $68.9$
@@ V-B Efficiency Analysis
To assess the computational efficiency of our proposed model, we conduct an efficiency analysis based on the results shown in Tab. II. The analysis reveals that our model not only achieves better performance but does so with significantly lower computational demands. For example, the model parameters are reduced by 81.2%, from 255.1 to 47.9 M, and memory usage decreased by 90.0% regarding the number of floating point operations (FLOPs). Additionally, we record the PyTorch allocated GPU memory during training and the required GPU usage dropped by 86.2%. Despite these reductions, our model improves the success rate (K=1) by 5.1% in Stick-Pull, 1.7% in Disassemble, and 3.3% in Pick-Out-Of-Hole. These results underscore the efficiency of our model, demonstrating its ability to deliver superior performance with much lower computational costs.
TABLE II: Efficiency Comparisons on Mamba Policy and 3D Diffusion Policy. We conduct experiments in difficult MetaWorld environments, where our method achieves stronger results with $\mathbf{80\%}$ fewer parameters. We abbreviate very hard and hard levels as VH and H.
Method # Param (M) FLOPs (G) GPU Usage (MB) Hard Environments in MetaWorld
Stick-Pull (VH) Disassemble (VH) Pick-Out-Of-Hole (H)
DP3 $255.1$ $0.30$ $996.1$ $81.6$ $80.0$ $45.0$
Mamba Policy $47.9$ ($\downarrow$ $\mathbf{81.2\%}$) $0.03$ ($\downarrow$ $\mathbf{90.0\%}$) $137.7$ ($\downarrow$ $\mathbf{86.2\%}$) $\mathbf{86.7}$ ($\uparrow$ $\mathbf{5.1\%}$) $\mathbf{81.7}$ ($\uparrow$ $\mathbf{1.7\%}$) $\mathbf{48.3}$ ($\uparrow$ $\mathbf{3.3\%}$)
@@ V-C Ablation Study
VI Conclusion
In this paper, we introduced the Mamba Policy, a lightweight yet efficient model tailored for 3D manipulation tasks. By significantly reducing the parameter count by over 80% compared to 3D diffusion policy, the Mamba Policy maintains strong performance while being more suitable for deployment on resource-constrained devices. Central to this approach is the XMamba Block, which effectively combines input data with conditional features through the integration of Mamba and Attention mechanisms. Our extensive experiments on the Adroit, Dexart, and MetaWorld datasets demonstrate the Mamba Policy’s superior performance and reduced computational requirements. Additionally, we showed that the Mamba Policy exhibits enhanced robustness in long-horizon scenarios, making it a compelling choice for various real-world applications. This work sets the stage for future research in developing efficient models for 3D manipulation, particularly in environments with limited computational resources. Our project page is in this link.

### v2

@@ Abstract
Diffusion models have been widely employed in the field of 3D manipulation due to their efficient capability to learn distributions, allowing for precise prediction of action trajectories. However, diffusion models typically rely on large parameter UNet backbones as policy networks, which can be challenging to deploy on resource-constrained devices. Recently, the Mamba model has emerged as a promising solution for efficient modeling, offering low computational complexity and strong performance in sequence modeling. In this work, we propose the Mamba Policy, a lighter but stronger policy that reduces the parameter count by over 80% compared to the original policy network while achieving superior performance. Specifically, we introduce the XMamba Block, which effectively integrates input information with conditional features and leverages a combination of Mamba and Attention mechanisms for deep feature extraction. Extensive experiments demonstrate that the Mamba Policy excels on the Adroit, Dexart, and MetaWorld datasets, requiring significantly fewer computational resources. Additionally, we highlight the Mamba Policy’s enhanced robustness in long-horizon scenarios compared to baseline methods and explore the performance of various Mamba variants within the Mamba Policy framework. Real-world experiments are also conducted to further validate its effectiveness. Our open-source project page can be found at https://andycao1125.github.io/mamba_policy/.
@@ IV-B XMamba
TABLE I: Quantitative comparisons of different baselines in simulation environments. We compare our Mamba Policy with IBC, BCRNN, 3D Diffusion Policy, and Diffusion Policy in Adroit, MetaWorld and DexArt datasets, in terms of SR5. denotes our reproduced results for fair comparison. Our proposed Mamba Policy achieves superior results across all domains.
@@
DP3 [4] $100.0\scriptstyle{\pm 0.0}$ $64.5\scriptstyle{\pm 2.4}$ $41.0\scriptstyle{\pm 3.3}$ $99.6\scriptstyle{\pm 0.4}$ $\mathbf{76.0\scriptstyle{\pm 2.4}}$ $\mathbf{100.0\scriptstyle{\pm 0.0}}$ $77.0\scriptstyle{\pm 3.5}$ $30.0\scriptstyle{\pm 2.9}$ $74.3\scriptstyle{\pm 2.6}$ $27.0\scriptstyle{\pm 0.8}$ $68.9$
@@ V Experiment
TABLE II: Efficiency comparisons on Mamba Policy and 3D Diffusion Policy. We conduct experiments in difficult MetaWorld environments, where our method achieves stronger results with $\mathbf{80\%}$ fewer parameters and $\mathbf{127\%}$ smaller GPU usage. We abbreviate very hard and hard levels as VH and H.
Method # Param (M) FLOPs (G) Torch Allocated Memory (MB) Total GPU Usage (MB) Hard Environments in MetaWorld
Stick-Pull (VH) Disassemble (VH) Pick-Out-Of-Hole (H)
DP3 $255.1$ $0.30$ $996.1$ $8306$ $81.6$ $80.0$ $45.0$
Mamba Policy $47.9$ ($\downarrow$ $\mathbf{81.2\%}$) $0.03$ ($\downarrow$ $\mathbf{90.0\%}$) $137.7$ ($\downarrow$ $\mathbf{86.2\%}$) $3648$ ($\downarrow$ $\mathbf{127.7\%}$) $\mathbf{86.7}$ ($\uparrow$ $\mathbf{5.1\%}$) $\mathbf{81.7}$ ($\uparrow$ $\mathbf{1.7\%}$) $\mathbf{48.3}$ ($\uparrow$ $\mathbf{3.3\%}$)
@@ V-B Efficiency Analysis
To assess the computational efficiency of our proposed model, we conduct an efficiency analysis based on the results shown in Tab. II. The analysis reveals that our model not only achieves better performance but does so with significantly lower computational demands. For example, under the Adorit Pen environment, the model parameters are reduced by 81.2%, from 255.1 to 47.9 M, and memory usage decreased by 90.0% regarding the number of floating point operations (FLOPs). Additionally, we record the PyTorch allocated GPU memory per loss backward during training and the required GPU usage dropped by 86.2%. The total GPU usage with our Mamba Policy is 127.7% smaller that the DP3’s. Despite these reductions, our model improves the success rate (K=1) by 5.1% in Stick-Pull, 1.7% in Disassemble, and 3.3% in Pick-Out-Of-Hole. These results underscore the efficiency of our model, demonstrating its ability to deliver superior performance with much lower computational costs.
@@ V-C Ablation Study
Figure 6: Real-world experiment. The robot successfully grasps the cup and places it on the cup holder as part of the 3D manipulation task. The sequence of images demonstrates the robot’s precise grasping and placement actions, validating the effectiveness of the proposed Mamba Policy in real-world robotic manipulation tasks.
VI Real World Experiment
Experiment Settings. In the real-world experiment, we utilize a Franka arm equipped with a gripper, along with a Realsense D455 camera to capture observations during the task. Inference is performed using a 4090 GPU. Our designed task involved the robot grasping a cup and placing it accurately on a cup holder. The robot is required to execute the task with high precision, both in terms of grasping the cup and positioning it correctly on the holder. Data collection is carried out via human teleoperation, where a total of 40 demonstrations are gathered for training the model.
Results. As illustrated in Fig. 6, our Mamba policy demonstrated a high level of precision in executing the grasping operation and successfully placing the cup on the holder. The results validate the effectiveness and practical applicability of the proposed method, showcasing the robot’s ability to perform the task with accuracy and reliability in a real-world environment.
VII Conclusion
In this paper, we introduced the Mamba Policy, a lightweight yet efficient model tailored for 3D manipulation tasks. By significantly reducing the parameter count by over 80% compared to 3D diffusion policy, the Mamba Policy maintains strong performance while being more suitable for deployment on resource-constrained devices. Central to this approach is the XMamba Block, which effectively combines input data with conditional features through the integration of Mamba and Attention mechanisms. Our extensive experiments on the Adroit, Dexart, and MetaWorld datasets demonstrate the Mamba Policy’s superior performance and reduced computational requirements. Additionally, we showed that the Mamba Policy exhibits enhanced robustness in long-horizon scenarios, making it a compelling choice for various real-world applications. This work sets the stage for future research in developing efficient models for 3D manipulation, particularly in environments with limited computational resources. Real-world experiments are also conducted to validate its effectiveness further. Our open-source project page can be found at this link.